OpenStudy (anonymous):

please help!! I know this simple but i keep messing up and i dont know where ((x^2/50)+20x+10)/x=(x/25)+20

OpenStudy (anonymous):

can you do the problem over for me so I can see where you went wrong?

OpenStudy (anonymous):

first i multiplied 50 out to the top and bottom getting (x^2+1000x+50)/50x=(x/25)+20

OpenStudy (littlebird):

It looks like the left half of the equation had everything over "x." Shouldn't you then have multiplied everything first by x?

OpenStudy (anonymous):

so i did what you said @Littlebird and i got (x^3/50)+20x^2+10x=(x/25)+20

OpenStudy (anonymous):

do you then multiply the whole thing by 50?

OpenStudy (anonymous):

@Preetha can you please help me?

OpenStudy (cwrw238):